Please kindly help me. I am having problem with this question.
The length of the swimming pool is 2x-1 meters, and the width is x+2 meters. Write a polynomial that represents the areaa. Find the area if x is 5 meters.
Thank you in advance.

Answer by checkley71(8403)

AREA=LENGTH*WIDTH
AREA=(2X-1)(X+2)
A=2X^2-X+4X-2
A=2X^2+3X-2 FORMULA FOR THE POOL AREA.
A=2*5^2+3*5-2
A=2*25+15-2
A=50+15-2
A=65-2
A=63 SQM.